# Stochastic Stability of ACIMs for Piecewise Expanding $C^{1+\varepsilon}$ Maps

### 04 Apr 2026 v1

@@ Abstract.
In the generalized bounded variation space $BV_{1,1/p}$, we establish a Lasota–Yorke inequality uniform in $\delta$. Consequently, each $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$, and $h_{\delta}\to h$ in $L^{1}$ as $\delta\to 0$, where $h$ is the ACIM density of $P_{\tau}$.
Our proof combines the $BV_{1,1/p}$ framework, adapted from recent ACIM existence results, with uniform quasi-compactness and perturbation theory for transfer operators. This establishes stochastic stability under minimal $C^{1+\varepsilon}$ regularity ($\varepsilon>0$), where the $C^{1}$ case is known to fail.
@@ 1. Introduction
Piecewise expanding maps of the interval form a fundamental class of dynamical systems in one dimension. A central object in their study is an absolutely continuous invariant measure (ACIM), whose density describes the long-term statistical behavior of almost all initial conditions. A classical result of Lasota and Yorke [1] establishes the existence of ACIMs for sufficiently smooth maps with uniform expansion via what is now known as the Lasota–Yorke inequality. This approach provides control over both variation and the $L^{1}$ norm of densities and leads to strong statistical properties such as mixing and decay of correlations. Subsequent works, including those of Jakobson [2] and Bowen [3], extended these results to broader settings, typically under higher regularity assumptions.
More recently, attention has turned to systems with minimal smoothness. In particular, [4] proved the existence of ACIMs for piecewise $C^{1+\varepsilon}$ expanding maps by working in the generalized bounded variation space $BV_{1,1/p}$ defined via oscillation seminorms:
(1.1) $\|f\|_{1,1/p}=\sup_{0<r\leq 1}\frac{|\mathrm{Osc}(f,r,\cdot)|_{1}}{r^{1/p}}+\|f\|_{1}.$
@@
A natural question is whether these invariant measures are stable under small perturbations. This problem, known as stochastic stability, concerns the persistence of invariant densities under the introduction of noise or small perturbations of the dynamics. It plays an important role in understanding the robustness of statistical properties of dynamical systems, particularly in applications where noise or numerical approximation is unavoidable. Classical results on stochastic stability have been established for uniformly expanding maps in spaces such as bounded variation or Hölder spaces, typically under stronger smoothness assumptions.
The purpose of this paper is to study stochastic stability of ACIMs for piecewise expanding $C^{1+\varepsilon}$ maps in the above low-regularity setting. Let $\tau$ be a map in the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps of the interval. We consider a family of perturbed Frobenius–Perron operators of the form
@@
where $P_{\tau}$ is the Frobenius–Perron operator associated with $\tau$, and $Q_{\delta}$ is a Markov smoothing operator representing noise of intensity $\delta>0$. Working in the space $BV_{1,1/p}$, we establish a Lasota–Yorke inequality that is uniform in $\delta$. As a consequence, each operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$, and we prove that $\|h_{\delta}-h\|_{L^{1}}\to 0$ as $\delta\to 0$, where $h$ is the invariant density of the unperturbed system $P_{\tau}$ in $BV_{1,1/p}$.
Our approach combines the generalized bounded variation framework with perturbation techniques for quasi-compact operators. This extends stochastic stability results to piecewise $C^{1+\varepsilon}$ maps under minimal regularity assumptions in the $BV_{1,1/p}$ setting.
The paper is organized as follows. In Section 2, we recall the definition and properties of the space $\mathrm{BV}^{1,1}_{p}$ and define the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps. In Section 3, we define the perturbation operators and establish a uniform Lasota–Yorke inequality. Section 4 proves existence of invariant densities and stochastic stability via convergence in $L^{1}$. Section 5 discusses remarks and further directions.
@@ 2. Preliminaries
In this section, we recall the definition and basic properties of the generalized bounded variation spaces $BV_{1,1/p}$, following ([4, 5]). These spaces provide the appropriate functional setting for studying the Frobenius–Perron operator associated with piecewise expanding maps of low regularity.
Let $I=[0,1]$, and let $m$ denote the Lebesgue measure on II. For a measurable function $f:I\to\mathbb{R}$ and $r>0$, define the oscillation of $f$ at scale $r$ by
(2.1) $\mathrm{Osc}(f,r,x)=\sup\bigl\{|f(y_{1})-f(y_{2})|:y_{1},y_{2}\in(x-r,x+r)\cap I\bigr\}.$
The function $\mathrm{Osc}(f,r,\cdot)$ is measurable and describes the local variation of $f$.
@@
Let $p\geq 1$. The space $BV_{1,1/p}$ consists of all functions $f\in L^{1}(I)$ such that
(2.3) $\mathrm{var}_{1,1/p}(f):=\sup_{0<r\leq 1}\frac{\mathrm{Osc}_{1}(f,r)}{r^{1/p}}<\infty.$
@@
(2.4) $\|f\|_{1,1/p}=\mathrm{var}_{1,1/p}(f)+\|f\|_{1}.$
The space $(BV_{1,1/p},\|\cdot\|_{1,1/p})$ is a Banach space and is continuously embedded in $L^{1}(I)$. Moreover, the unit ball in $BV_{1,1/p}$ is relatively compact in $L^{1}(I)$.
@@
    $\mathrm{Osc}(f,r,\cdot)$ is lower semicontinuous and hence measurable;
@@
(2.5) $\mathrm{Osc}(f\circ\tau^{-1},r,y)\leq\mathrm{Osc}\left(f,\frac{r}{s},\tau^{-1}(y)\right)$
@@
Assumption 2.5 (Expansion condition (q3)).
By Theorem 3 of [4], for $\tau\in\mathcal{T}([0,1];s,\varepsilon)$, the Frobenius–Perron operator satisfies the Lasota–Yorke inequality
$\mathrm{var}_{1,1/p}(P_{\tau}f)\leq\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}$
with contraction constant $\alpha_{0}=s^{-1/p}+2s^{-1}$ and $C_{0}$ depending only on $\tau$. The hypothesis $\alpha_{0}<1/C_{1}$ of Theorem 3.5 is therefore equivalent to
$\frac{1}{s^{1/p}}+\frac{2}{s}<\frac{1}{2^{1/p}\max(1,4\|q^{\prime}\|_{\infty})},$
which holds for all sufficiently large $s$. We assume henceforth that $s$ satisfies this stronger condition.
@@ 3. Perturbation Operators and Uniform Lasota–Yorke Inequality
Let $q:\mathbb{R}\to[0,\infty)$ be a $C^{1}$ function satisfying
(3.1) $\int_{\mathbb{R}}q(z)\,dz=1,\qquad\mathrm{supp}(q)\subset[-1,1].$
For $\delta\in(0,1/4)$, define the kernel $q_{\delta}(x,y)=\delta^{-1}q((x-y)/\delta)$ and the smoothing operator
$(Q_{\delta}f)(x)\;=\;\int_{I}q_{\delta}(x,y)\,f(y)\,dy,\qquad x\in I.$
Because $\mathrm{supp}(q)\subset[-1,1]$, the kernel $q_{\delta}(x,\cdot)$ is supported on $(x-\delta,\,x+\delta)$. The perturbed operator is $P_{\delta}=Q_{\delta}P_{\tau}$.
For the oscillation estimates in Lemmas 3.3 and 3.4 we use the shift-invariance of the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. Specifically, for $x\in[\delta,\,1-\delta]$ the support of $q_{\delta}(x,\cdot)$ lies entirely within $(0,1)\subset I$, so the torus distance $d_{\mathbb{T}}$ and the Euclidean distance agree on this support, and the torus computation is identical to the interval computation. For $x$ in the boundary strips $0,\delta)\cup(1-\delta,1]$, the kernel reaches outside II; these regions contribute at most $C\|g\|_{1}$ to $\mathrm{Osc}_{1}$ and are handled by Lemma 3.2. All subsequent oscillation estimates on II follow by combining the interior torus bound with Lemma 3.2.
@@
$\int P_{\delta}f\,dm=\int f\,dm\quad\text{for all }f\in L^{1}(I),$
and is a contraction on $L^{1}(I)$:
$\|P_{\delta}f\|_{1}\leq\|f\|_{1}\quad\text{for all }f\in L^{1}(I).$
Moreover, if $f\geq 0$, then $\|P_{\delta}f\|_{1}=\|f\|_{1}$.
@@
The operators $P_{\tau}$ and $Q_{\delta}$ are positive linear operators. Moreover, $P_{\tau}$ preserves integrals, and for $Q_{\delta}$ we have
$(Q_{\delta}f)(x)=\int_{\mathbb{T}}q_{\delta}(x,y)f(y)\,dy,$
where $q_{\delta}(x,y)\geq 0$ and $\int_{\mathbb{T}}q_{\delta}(x,y)\,dx=1$ for each $y\in\mathbb{T}$ (since $\int_{\mathbb{R}}q_{\delta}(x,y)\,dx=1$ and wraparound does not occur). Thus,
$\int_{\mathbb{T}}(Q_{\delta}f)(x)\,dx=\int_{\mathbb{T}}f(y)\left[\int_{\mathbb{T}}q_{\delta}(x,y)\,dx\right]dy=\int_{\mathbb{T}}f(y)\,dy=\int_{I}f(y)\,dy.$
Thus,
$\int(Q_{\delta}f)(x)\,dx=\int f(y)\,dy,$
so $Q_{\delta}$ also preserves integrals. Hence $P_{\delta}=Q_{\delta}P_{\tau}$ is positive and preserves integrals.
To prove the $L^{1}$ contraction property, note that for any $f\in L^{1}(I)$,
$|P_{\delta}f|\leq P_{\delta}|f|,$
by positivity. Integrating and using preservation of integrals, we obtain
$\|P_{\delta}f\|_{1}=\int|P_{\delta}f|\,dm\leq\int P_{\delta}|f|\,dm=\int|f|\,dm=\|f\|_{1}.$
Finally, if $f\geq 0$, then $|P_{\delta}f|=P_{\delta}f$, so
$\|P_{\delta}f\|_{1}=\int P_{\delta}f\,dm=\int f\,dm=\|f\|_{1}.$
Lemma 3.2 (Boundary strip estimate).
For any $g\in L^{1}(I)$, $r>0$, and $0<\delta<1/4$,
$\int_{0}^{\delta}\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\,dx\;+\;\int_{1-\delta}^{1}\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\,dx\;\leq\;4\|q\|_{\infty}\|g\|_{1}.$
Proof.
For any $x\in I$, pointwise:
$\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\;\leq\;2\|Q_{\delta}g\|_{\infty}\;\leq\;\frac{2\|q\|_{\infty}}{\delta}\|g\|_{1},$
where the last bound uses $|(Q_{\delta}g)(x)|\leq\|q_{\delta}(x,\cdot)\|_{\infty}\|g\|_{1}=\delta^{-1}\|q\|_{\infty}\|g\|_{1}$. Integrating over $0,\delta)$ and $(1-\delta,1]$, each of measure $\delta$:
$\int_{0}^{\delta}\mathrm{Osc}(Q_{\delta}g,r,x)\,dx\;\leq\;\delta\cdot\frac{2\|q\|_{\infty}}{\delta}\|g\|_{1}\;=\;2\|q\|_{\infty}\|g\|_{1}.$
Summing both boundary strips gives the result. ∎
Lemma $3.$
Work on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ with $0<\delta<1/2$. Let $q_{\delta}(x,y)=\frac{1}{\delta}q\!\left(\frac{x-y}{\delta}\right)$ where $q\in C^{1}$, $\operatorname{supp}(q)\subset[-1,1]$, $q\geq 0$, $\int q=1$. Set $C:=4\|q^{\prime}\|_{\infty}$. For all $g\in L^{1}(\mathbb{T})$ and $0<r\leq\delta$,
$\mathrm{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\frac{Cr}{\delta}\,\mathrm{Osc}_{1}(g,\,r+\delta).$
@@
Set $c:=\dfrac{1}{|B(x,r+\delta)|}\displaystyle\int_{B(x,r+\delta)}g(y)\,dy$. Since $\int q_{\delta}(x_{i},y)\,dy=1$ for each $i$,
@@
The integrand is supported inside $B(x,r+\delta)$. For $y\in B(x,r+\delta)$, since both $y$ and the averaging variable $z$ lie in $B(x,r+\delta)$,
$|g(y)-c|\;=\;\left|\frac{1}{|B(x,r+\delta)|}\int_{B(x,r+\delta)}\bigl(g(y)-g(z)\bigr)\,dz\right|\;\leq\;\sup_{z\in B(x,r+\delta)}|g(y)-g(z)|\;\leq\;\mathrm{Osc}(g,\,r+\delta,\,x).$
For the kernel difference, substitute $u=(x_{1}-y)/\delta$ and apply the mean value theorem; since $\operatorname{supp}(q)\subset[-1,1]$,
$\int_{\mathbb{T}}\bigl|q_{\delta}(x_{1},y)-q_{\delta}(x_{2},y)\bigr|\,dy=\int_{\mathbb{R}}\left|q(u)-q\!\Bigl(u-\tfrac{x_{1}-x_{2}}{\delta}\Bigr)\right|du\;\leq\|q^{\prime}\|_{L^{1}}\cdot\frac{|x_{1}-x_{2}|}{\delta}\leq 2\|q^{\prime}\|_{\infty}\cdot\frac{2r}{\delta}=\frac{Cr}{\delta}$
where $\|q^{\prime}\|_{L^{1}}\leq|\operatorname{supp}(q^{\prime})|\cdot\|q^{\prime}\|_{\infty}\leq 2\|q^{\prime}\|_{\infty}$ (since $\operatorname{supp}(q)\subset[-1,1]$), and $|x_{1}-x_{2}|\leq 2r$. Combining, $\mathrm{Osc}(Q_{\delta}g,\,r,\,x)\leq\dfrac{Cr}{\delta}\,\mathrm{Osc}(g,\,r+\delta,\,x)$. Integrating over $x\in\mathbb{T}$ gives the result. ∎
Lemma 3.4 (Translation coupling).
@@
$\mathrm{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\mathrm{Osc}_{1}(g,\,r+\delta).$
@@
Fix $x\in\mathbb{T}$ and $x_{1},x_{2}\in B(x,r)$. Set $v:=x_{2}-x_{1}$. Since $q_{\delta}(x_{2},y)=q_{\delta}(x_{1},y-v)$, the substitution u=yvu=y-v on $\mathbb{T}$ gives
$(Q_{\delta}g)(x_{2})=\int_{\mathbb{T}}q_{\delta}(x_{1},u)\,g(u+v)\,du,$
@@
The kernel $q_{\delta}(x_{1},\cdot)$ is supported on $x_{1}-\delta,\,x_{1}+\delta]\subset B(x,\,r+\delta)$. For each $u$ in this support: $u\in[x_{1}-\delta,\,x_{1}+\delta]\subset B(x,\,r+\delta)$, and $u+v\in[x_{2}-\delta,\,x_{2}+\delta]\subset B(x,\,r+\delta)$  (since $|(u+v)-x_{2}|=|u-x_{1}|\leq\delta$). Hence $|g(u)-g(u+v)|\leq\mathrm{Osc}(g,\,r+\delta,\,x)$. Since $\int q_{\delta}(x_{1},u)\,du=1$,
$|(Q_{\delta}g)(x_{1})-(Q_{\delta}g)(x_{2})|\;\leq\;\mathrm{Osc}(g,\,r+\delta,\,x).$

### 11 Apr 2026 v2

@@ Abstract.
In the generalized bounded variation space $BV_{1,1/p}(I)$, we establish a Lasota–Yorke inequality uniform in $\delta$. Consequently, each $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}(I)$, and $h_{\delta}\to h$ in $L^{1}(I)$ as $\delta\to 0$, where $h$ is the ACIM density of $P_{\tau}$.
Our proof combines the $BV_{1,1/p}(I)$ framework with uniform quasi-compactness and the Keller–Liverani perturbation theory for transfer operators. Under a topological mixing assumption on $\tau$, this establishes stochastic stability at the $C^{1+\varepsilon}$ threshold, which is essentially optimal for ACIM existence: $C^{1}$ regularity alone is insufficient to guarantee the existence of an ACIM.
@@ 1. Introduction
Piecewise expanding maps of the interval form a fundamental class of dynamical systems in one dimension. A central object in their study is an absolutely continuous invariant measure (ACIM), whose density describes the long-term statistical behavior of almost all initial conditions. A classical result of Lasota and Yorke [1] establishes the existence of ACIMs for sufficiently smooth maps with uniform expansion via what is now known as the Lasota–Yorke inequality. This approach provides control over both variation and the $L^{1}(I)$ norm of densities and leads to strong statistical properties such as mixing and decay of correlations. Subsequent works by Pianigiani-Yorke [7], Hofbauer–Keller [8], and Keller [3] extended these results to broader settings, typically under higher regularity assumptions.
More recently, attention has turned to systems with minimal smoothness. In particular, Rajput and Góra [2] proved the existence of ACIMs for piecewise $C^{1+\varepsilon}$ expanding maps by working in the generalized bounded variation space $BV_{1,1/p}(I)$ defined via oscillation seminorms:
(1.1) $\|f\|_{1,1/p}=\sup_{0<r\leq 1}\frac{\operatorname{Osc}_{1}(f,r)}{r^{1/p}}+\|f\|_{1}.$
@@
A natural question is whether these invariant measures are stable under small perturbations. This problem, known as stochastic stability, concerns the persistence of invariant densities under the introduction of noise or small perturbations of the dynamics. It plays an important role in understanding the robustness of statistical properties of dynamical systems, particularly in applications where noise or numerical approximation is unavoidable. Classical results on stochastic stability have been established for uniformly expanding maps in spaces such as bounded variation or Hölder spaces, typically under stronger smoothness assumptions. In the piecewise $C^{1+\varepsilon}$ setting specifically, Góra [9] proved stochastic stability using Rychlik’s $C_{\varepsilon}$ seminorm, under expansion conditions more restrictive than those required here. The present paper reproves and strengthens this result in the $BV_{1,1/p}(I)$ framework via the Keller–Liverani perturbation theory, with weaker hypotheses and a more transparent spectral argument.
The purpose of this paper is to study stochastic stability of ACIMs for piecewise expanding $C^{1+\varepsilon}$ maps in the above low-regularity setting. Let $\tau$ be a map in the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps of the interval. The noise operator $Q_{\delta}$ is defined via a periodized (torus) convolution kernel, which automatically satisfies the Markov property
$\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,dy=1\quad\text{for all }x\in I,$
without requiring a separate boundary correction. This is the mechanism that allows us to work directly on the interval $I=[0,1]$. We consider a family of perturbed Frobenius–Perron operators of the form
@@
where $P_{\tau}$ is the Frobenius–Perron operator associated with $\tau$, and $Q_{\delta}$ is a Markov smoothing operator representing noise of intensity $\delta>0$. Working in the space $BV_{1,1/p}(I)$, we establish a Lasota–Yorke inequality that is uniform in $\delta$. As a consequence, each operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}(I)$, and we prove that $\|h_{\delta}-h\|_{L^{1}(I)}\to 0$ as $\delta\to 0$, where $h$ is the invariant density of the unperturbed system $P_{\tau}$ in $BV_{1,1/p}(I)$.
Our approach combines the generalized bounded variation framework with perturbation techniques for quasi-compact operators. This extends stochastic stability results to piecewise $C^{1+\varepsilon}$ maps under minimal regularity assumptions in the $BV_{1,1/p}(I)$ setting.
The paper is organized as follows. In Section 2, we recall the definition and properties of the space $BV_{1,1/p}(I)$ and define the class $\mathcal{T}([0,1];s,\varepsilon)$ of piecewise expanding $C^{1+\varepsilon}$ maps. In Section 3, we define the perturbation operators and establish a uniform Lasota–Yorke inequality. Section 4 proves existence of invariant densities and stochastic stability via convergence in $L^{1}(I)$. Section 5 presents a nonlinear example illustrating the theory. Section 6 discusses remarks and further directions.
@@ 2. Preliminaries
In this section, we recall the definition and basic properties of the generalized bounded variation spaces $BV_{1,1/p}(I)$, following [2, 3]. These spaces provide the appropriate functional setting for studying the Frobenius–Perron operator associated with piecewise expanding maps of low regularity.
Let $I=[0,1]$, and let $m$ denote the Lebesgue measure on $I$. For a measurable function $f:I\to\mathbb{R}$ and $r>0$, define the oscillation of $f$ at scale $r$ by
(2.1) $\operatorname{Osc}(f,r,x)=\sup\bigl\{|f(y_{1})-f(y_{2})|:y_{1},y_{2}\in(x-r,x+r)\cap I\bigr\}.$
The function $\operatorname{Osc}(f,r,\cdot)$ is measurable and describes the local variation of $f$.
@@
Let $p\geq 1$. The space $BV_{1,1/p}(I)$ consists of all functions $f\in L^{1}(I)$ such that
(2.3) $\operatorname{var}_{1,1/p}(f):=\sup_{0<r\leq 1}\frac{\operatorname{Osc}_{1}(f,r)}{r^{1/p}}<\infty.$
@@
(2.4) $\|f\|_{1,1/p}=\operatorname{var}_{1,1/p}(f)+\|f\|_{1}.$
The space $\bigl(BV_{1,1/p}(I),\|\cdot\|_{1,1/p}\bigr)$ is a Banach space and is continuously embedded in $L^{1}(I)$. Moreover, the unit ball in $BV_{1,1/p}(I)$ is relatively compact in $L^{1}(I)$.
@@
    $\operatorname{Osc}(f,r,\cdot)$ is lower semicontinuous and hence measurable;
@@
(2.5) $\operatorname{Osc}(f\circ\tau^{-1},r,y)\leq\operatorname{Osc}\left(f,\frac{r}{s},\tau^{-1}(y)\right)$
@@
Assumption 2.5.
[Lasota–Yorke condition for $P_{\tau}$] By Theorem 3.1 of [2], there exist constants $0<\alpha_{0}<1$ and $C_{0}>0$, depending only on $\tau$, such that
$\operatorname{var}_{1,1/p}(P_{\tau}f)\leq\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}\qquad\text{for all }f\in BV_{1,1/p}(I).$
Moreover, $\alpha_{0}\leq s^{-1/p}$ by [2, Theorem 3.1].
@@ 3. Perturbation Operators and Uniform Lasota–Yorke Inequality
Let $q:\mathbb{R}\to[0,\infty)$ be a $C^{1}$ function with compact support satisfying
$\int_{\mathbb{R}}q(z)\,dz=1,\qquad\operatorname{supp}(q)\subset[-1,1],$
and assume in addition that $q^{\prime}\in L^{\infty}(\mathbb{R})$. For $\delta\in(0,1/4)$, define the periodized kernel
$q_{\delta}^{\mathrm{per}}(x,y)=\sum_{n\in\mathbb{Z}}\delta^{-1}\,q\!\left(\frac{x-y+n}{\delta}\right),\qquad x,y\in I,$
and the smoothing operator
$(Q_{\delta}f)(x)=\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,f(y)\,dy,\qquad x\in I.$
Assumption 3.1.
We assume that $\tau$ and $q$ satisfy the joint expansion–kernel condition
$\alpha_{0}\;<\;\frac{1}{\widetilde{C}_{1}},\qquad\widetilde{C}_{1}:=(1+2^{1+1/p})\cdot 2^{1/p}\max\!\bigl\{1,\,4\|q^{\prime}\|_{\infty}\bigr\}.$
Since $\alpha_{0}\leq s^{-1/p}$ (Assumption 2.5), this holds whenever $s^{1/p}>\widetilde{C}_{1}$, i.e., $s>\widetilde{C}_{1}^{\,p}$. For any fixed $\tau$, it is satisfied by choosing $q$ with $\|q^{\prime}\|_{\infty}$ small enough that $4\|q^{\prime}\|_{\infty}\leq 1$ and $s^{1/p}>(1+2^{1+1/p})\cdot 2^{1/p}$.
Since $\operatorname{supp}(q)\subset[-1,1]$ and $\delta<1/4$, for each pair $(x,y)\in I\times I$ at most two terms in the sum are nonzero. The substitution $u=(x-y+n)/\delta$ gives
$\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,dx=\sum_{n\in\mathbb{Z}}\int_{(n-y)/\delta}^{(n+1-y)/\delta}q(u)\,du=\int_{-\infty}^{\infty}q(u)\,du=1\quad\text{for all }y\in I,$
and the identical argument with the roles of $x$ and $y$ interchanged gives $\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,dy=1$ for all $x\in I$, so the kernel is doubly stochastic on $I$. For $x\in[\delta,1-\delta]$ the support of $q_{\delta}^{\mathrm{per}}(x,\cdot)$ lies entirely inside $I$, so $q_{\delta}^{\mathrm{per}}(x,y)=\delta^{-1}q((x-y)/\delta)$ there; the periodization affects only the strips $0,\delta)\cup(1-\delta,1]$. Since $q_{\delta}^{\rm per}$ is precisely the convolution kernel of $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ restricted to $I\times I$, the operator $Q_{\delta}$ on $I$ coincides with torus convolution. We identify functions on $I$ with their periodic extensions to $\mathbb{T}$ when applying convolution operators. The perturbed operator is $P_{\delta}=Q_{\delta}P_{\tau}$. The oscillation estimates established in Lemmas 3.3 and  3.4 below therefore apply directly to $Q_{\delta}$ on $I$ without any boundary correction.
@@
$\int_{I}P_{\delta}f\,dm=\int_{I}f\,dm\quad\text{for all }f\in L^{1}(I),$
and is a contraction on $L^{1}(I)$: $\|P_{\delta}f\|_{1}\leq\|f\|_{1}$ for all $f\in L^{1}(I)$. Moreover, if $f\geq 0$ then $\|P_{\delta}f\|_{1}=\|f\|_{1}$.
@@
The operators $P_{\tau}$ and $Q_{\delta}$ are positive and linear. The Frobenius–Perron operator $P_{\tau}$ preserves integrals by construction.
For $Q_{\delta}$, since $q_{\delta}^{\mathrm{per}}\geq 0$, Tonelli’s theorem gives
$\int_{I}\!\int_{I}q_{\delta}^{\mathrm{per}}(x,y)|f(y)|\,dm(y)\,dm(x)=\int_{I}|f(y)|\underbrace{\left[\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,dm(x)\right]}_{=\,1}dm(y)=\|f\|_{1}<\infty,$
so the double integral is absolutely convergent. Fubini’s theorem then permits interchange of integration order for $f$ itself:
$\int_{I}(Q_{\delta}f)(x)\,dm(x)=\int_{I}f(y)\left[\int_{I}q_{\delta}^{\mathrm{per}}(x,y)\,dm(x)\right]dm(y)=\int_{I}f(y)\,dm(y).$
Hence $Q_{\delta}$ preserves integrals, and so does $P_{\delta}=Q_{\delta}P_{\tau}$.
For the $L^{1}(I)$ bound, positivity of $P_{\delta}$ and linearity give $|P_{\delta}f|\leq P_{\delta}|f|$ pointwise. Integrating and applying integral preservation (with $|f|\geq 0$):
$\|P_{\delta}f\|_{1}=\int_{I}|P_{\delta}f|\,dm\leq\int_{I}P_{\delta}|f|\,dm=\int_{I}|f|\,dm=\|f\|_{1}.$
If $f\geq 0$, then $|P_{\delta}f|=P_{\delta}f$, so equality holds. Thus $Q_{\delta}$, and hence $P_{\delta}$, is a Markov operator. ∎
Lemma 3.3 (Small-scale oscillation on $\mathbb{T}$).
Work on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ with $0<\delta<1/2$. Let $q_{\delta}(x,y)=\tfrac{1}{\delta}q\!\left(\tfrac{x-y}{\delta}\right)$, where $q\in C^{1}$, $\operatorname{supp}(q)\subset[-1,1]$, $q\geq 0$, $\int_{\mathbb{R}}q(t)\,dt=1$. Set $C:=4\|q^{\prime}\|_{\infty}$. For all $g\in L^{1}(\mathbb{T})$ and $0<r\leq\delta$,
$\operatorname{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\frac{Cr}{\delta}\,\operatorname{Osc}_{1}(g,\,r+\delta).$
@@
Set $c:=\frac{1}{|B(x,r+\delta)|}\int_{B(x,r+\delta)}g(y)\,dy$. Since $\int q_{\delta}(x_{i},y)\,dy=1$ for each $i$,
@@
The integrand is supported inside $B(x,r+\delta)$. For $y\in B(x,r+\delta)$, $|g(y)-c|\leq\operatorname{Osc}(g,r+\delta,x)$. For the kernel difference, substituting $u=(x_{1}-y)/\delta$ and using the mean value theorem and the fact that $\operatorname{supp}(q^{\prime})\subset[-1,1]$, so that $\|q^{\prime}\|_{1}\leq 2\|q^{\prime}\|_{\infty}$, we obtain
$\int_{\mathbb{T}}|q_{\delta}(x_{1},y)-q_{\delta}(x_{2},y)|\,dy\leq 2\|q^{\prime}\|_{\infty}\cdot\frac{|x_{1}-x_{2}|}{\delta}\leq\frac{Cr}{\delta},$
where $C=4\|q^{\prime}\|_{\infty}$ and $|x_{1}-x_{2}|\leq 2r$. Combining, $\operatorname{Osc}(Q_{\delta}g,r,x)\leq\frac{Cr}{\delta}\operatorname{Osc}(g,r+\delta,x)$. Integrating over $x\in\mathbb{T}$ gives the result. ∎
Lemma 3.4 (Translation coupling on $\mathbb{T}$).
@@
$\operatorname{Osc}_{1}(Q_{\delta}g,\,r)\;\leq\;\operatorname{Osc}_{1}(g,\,r+\delta).$
@@
Fix $x\in\mathbb{T}$ and $x_{1},x_{2}\in B(x,r)$. Set $v:=x_{2}-x_{1}$. Since $q_{\delta}(x_{2},y)=q_{\delta}(x_{1},y-v)$, by translation invariance on $\mathbb{T}$, the change of variables u=yvu=y-v gives
$(Q_{\delta}g)(x_{2})=\int_{\mathbb{T}}q_{\delta}(x_{1},u)\,g(u+v)\,du.$
@@
The kernel $q_{\delta}(x_{1},\cdot)$ is supported on $x_{1}-\delta,x_{1}+\delta]\subset B(x,r+\delta)$. For each $u$ in this support: $u\in B(x,r+\delta)$, and $u+v\in[x_{2}-\delta,x_{2}+\delta]\subset B(x,r+\delta)$ (since $|(u+v)-x_{2}|=|u-x_{1}|\leq\delta$). Hence $|g(u)-g(u+v)|\leq\operatorname{Osc}(g,r+\delta,x)$. Since $\int q_{\delta}(x_{1},u)\,du=1$,
$|(Q_{\delta}g)(x_{1})-(Q_{\delta}g)(x_{2})|\leq\operatorname{Osc}(g,r+\delta,x).$